OpenStudy (anonymous):

there is a lock ..only 4 digits are required to open it.if the numbers are (0-9) you can use the numbers more then once...In how many ways are there to open the lock? PLEASE HELP MEDAL!!!

OpenStudy (danjs):

0000 to 9999

OpenStudy (anonymous):

9999?

OpenStudy (danjs):

+1

OpenStudy (danjs):

include 0000

OpenStudy (anonymous):

yeah the numbers are 0 to 9 and they can repeat

OpenStudy (anonymous):

Is this a permutation or combination not sure

OpenStudy (danjs):

there are 10 numbers to choose from (0,1,2...,9)

OpenStudy (anonymous):

thanks for that site I will definantly check it out...............yeah there are ten number

OpenStudy (anonymous):

then what the next step would be?

OpenStudy (anonymous):

The problem is a permutations with repetition... 10^4 = 10 000 ways to open

OpenStudy (danjs):

look at the permutations with repetition , it is towards the top of that page

OpenStudy (danjs):

the number of choices for each place, does not reduce if a number is already used, so you have 10 choices for the first space, 10 for the second... and so on

OpenStudy (anonymous):

ok I got if....but now what if the number cannot repeat..??

OpenStudy (anonymous):

I just need to multiply 9x8x7x6??

OpenStudy (danjs):

then if a number is used , the next space has 1 less to choose from

OpenStudy (danjs):

that is where the factorial thing comes in

OpenStudy (danjs):

but, you want to cut it off depending on how many numbers you are choosing , here is 4

OpenStudy (danjs):

so take 10 factorial, then divide that by (10-4) factorial

OpenStudy (danjs):

that will give you \[\frac{ 10*9*8*7*6*5*4*3*2*1 }{ 6*5*4*3*2*1 } = 10*9*8*7\]

OpenStudy (anonymous):

5040

OpenStudy (danjs):

yeah, if you cant have double numbers in the combination

OpenStudy (anonymous):

alright.thanks a lot for you help DanJs

OpenStudy (danjs):

welcome, use that page i linked, it is all there pretty clear.